OpenStudy (cinniomon):

Use the derivitive to identify the open intervals on which the function is increasing or decreasing. f(x)=x^3/4 - 3x

OpenStudy (australopithecus):

first off take the derivative

OpenStudy (cinniomon):

Ok

OpenStudy (cinniomon):

Then what

OpenStudy (australopithecus):

Then check to see where the derivative is 0, and then check the derivatives domain

OpenStudy (australopithecus):

to get the critical numbers

OpenStudy (australopithecus):

If you get a critical number that isn't in the domain of the original function do not use it

OpenStudy (australopithecus):

Give me the critical numbers you get it is always smart to write on a test set f'(x) = 0

OpenStudy (australopithecus):

that is worth a mark or part of a mark at least because it shows you know what you are doing

OpenStudy (australopithecus):

Once you give me the critical numbers I will show you how to figure out in what intervals your function is increasing and decreasing

OpenStudy (cinniomon):

Ok is one of the numbers the sq root of 8/3?

OpenStudy (australopithecus):

ok you only got one critical number?

OpenStudy (australopithecus):

I'm studying for an exam while helping you so you will have to do most of the work if that is alright

OpenStudy (australopithecus):

I can show you the method of solving it though

OpenStudy (cinniomon):

That would be great if you could show me the method and its fine if you r studying

OpenStudy (australopithecus):

just say you only have one critical number Make a table like so:  Now plug in any number in the interal \[(-\infty, \sqrt{\frac{8}{3}})\] into f(x) and see if you get a positive or negative, the number doesn't matter just the sign Do the same with this interval \[(\sqrt{\frac{8}{3}},\infty)\] Tell me what your results are and I will show you how to fill out the table